OpenStudy (anonymous):

A spring is fixed at the bottom end of an incline of inclination 37 degrees. A small block of mass 2 kg starts slipping down the incline from a point 4.8m away from the spring. The block compresses the spring by 20 cm, stops momentarily and then rebounds through a distance 1 m up the incline. Find (a) the friction coefficient between block and plane and (b) The spring constant 'k'

OpenStudy (missmob):

4.8mgsin(35°) - 4.8μmgsin(35°) = (1/2)(0.2)^2k + 0.2μmgsin(35°) (1/2)(0.2)^2k - 0.2μmgsin(35°) = mgsin(35°) + μmgsin(35°) 4.8mgsin(35°) - 5μmgsin(35°) = (1/2)(0.2)^2k (1/2)(0.2)^2k - 1.2μmgsin(35°) = mgsin(35°) mgsin(35°)(4.8 - 5μ) = (1/2)(0.2)^2k (1/2)(0.2)^2k = mgsin(35°)(1 + 1.2μ) 4.8 - 5μ = 1 + 1.2μ 6.2μ = 3.8 μ = 0.6129 k = 976.19 N/m

OpenStudy (anonymous):

can u please explain

OpenStudy (missmob):

im not good at explanations

OpenStudy (anonymous):

i am not being able to make anything from the answer.Any kind of explaation will be helpful

OpenStudy (missmob):

ahh?

OpenStudy (anonymous):

please!!!

OpenStudy (missmob):

;(

OpenStudy (yrelhan4):

ok you need to apply conservation of energy here.. you should know what that is, and how to apply that.. yeah?

OpenStudy (anonymous):

ya , i am getting confused in the second case when is rebounds. for the first case i have made the equation : k/25 = 60 - 40u k - spring constant and u - coefficient of ffriction

OpenStudy (yrelhan4):

can you write the full initial equation? should be 4.8mgsin(35°) - 4.8μmgsin(35°) = (1/2)(0.2)^2k + 0.2μmgsin(35°), as missMob said..

OpenStudy (yrelhan4):

for the first case i mean. does that match?

OpenStudy (missmob):

thats wat i did

OpenStudy (yrelhan4):

lol. you couldnt explain it though. but yeah. well done. :)

OpenStudy (missmob):

thx

OpenStudy (anonymous):

no the total distance it will travel will be 5 mtr and height will be 3 mtrs(using sin37) so spring energy = mgh - w.d by friction k/25 = 2X10X3 - 2*!0*cos37 u *5

OpenStudy (anonymous):

we can get 5 mtr as 4.8 + 20cm = 5 mte

OpenStudy (dls):

yup

OpenStudy (missmob):

sorry i couldnt explain im really bad at explinations lol :/

OpenStudy (yrelhan4):

well i guess cons of energy should be used here. i kinda understand what you did there, but i'll still stick with cons of energy.

OpenStudy (anonymous):

i used that only: innital energy = mgh + energy lost due to friction final energy = 1/2kx^2

OpenStudy (yrelhan4):

forgot friction?

OpenStudy (anonymous):

no

OpenStudy (yrelhan4):

oh i see it now.. the first equation looks right. for the rebound, initial energy of the spring is 1/2kx^2.. final energy should be 1.02*(mu)*mgcos35 + mgsin(35)..

OpenStudy (yrelhan4):

.02*mu*mgcos35 for the part when it rebounds to its equilibrium position.. and mu*mg*mgcos35 for the part where it goes up another 1m up the incline..

OpenStudy (yrelhan4):

yeah?

OpenStudy (anonymous):

ya i think its right

OpenStudy (yrelhan4):

yeah. now two variables, two equations.. solve them to get your answer.